OpenStudy (anonymous):

Integrate sin3xcos3x dx

OpenStudy (anonymous):

will give a medal please need help

OpenStudy (anonymous):

\[\int\limits \sin3xcos3x= \int\limits \frac{1}{2} \times 2\sin3xcos3x= \frac{1}{2} \int\limits \sin6x\] Let 6x=t Therefore: 6dx=dt i.e. dx=1/6dt Hence \[ \frac{1}{2} \int\limits \sin6x = \frac{1}{2} \int\limits sint \frac{dt}{6} = \frac{1}{2} \times \int\limits sint \frac{dt}{6} \] \[= \frac{1}{2} \times \frac{1}{6} (-cost)+c = -\frac{1}{12}cost+c\] \[= -\frac{1}{12}\cos6x+c\] (Since t=6x) @07supatel

OpenStudy (anonymous):

why is there 1/2 there

OpenStudy (anonymous):

thats the bit i dont understand

OpenStudy (anonymous):

@07supatel We know that: \[\sin2 \theta = 2\sin \theta \cos \theta \rightarrow \sin2 \theta = 2\sin \theta \cos \theta\] From the formula it is clear that: In order to write sinθ cosθ as sin2θ we should have written 2sinθcosθ =sin2θ i.e. 2sinθcosθ =sin2θ or vice versa similarly \[\sin3xcos3x = \frac{1}{2} \times 2 \sin3xcos3x = \frac{1}{2} \times \sin(2 \times3x)\] \[= \frac{1}{2} \sin(6x)\] Here 2 is multiplied and divided to balance the ratio.